# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ A.4.1  Assumptions
  • Choice of other graph input parameters: We saw earlier the effect of edge weight choice on the prospect of realizing an advantage. Other graph input parameters such as choice of seed value may also influence our conclusions. For our analyses presented in the main text, we fix the seed value for reproducibility of our results. To understand the influence of seed values, we pick three representative candidates: Barabási-Albert graph family from Laplacian matrix based system of linear equations, Gaussian random partition graph (no source and sink) family, and planted partition graph (no source and sink) family from incidence matrix based system of linear equations. The results of our analyses are presented in Fig. A.7. For each graph family, we pick three different seed values and analyze the advantage offered by the respective graph families. We find that for the candidates that we consider, there is no change in their category upon changing the seed values. For example, the Barabási-Albert graph, which is a good graph, remains a good graph. Although the categories do not change for the graph families for the seed values that we considered for them, our observations cannot be generalized, and this aspect requires further analysis in a future study. The effect of tuning other parameters on advantage is also beyond the scope of our current work, and we defer it for a future study.
A.4.2  Scope of our numerical study
  • Number of quantum linear solvers considered: Although we consider several efficient QLSs for our survey, we do exclude some, such as the quantum singular value transformation (QSVT)-based QLS. While QSVT is important due to its potential to offer near-optimal scaling in $\kappa$ and $\epsilon$, we do not consider it due to the non-trivial nature of $\mathcal{N}$ dependence entering via block encoding costs in its complexity. Other near-optimally scaling approaches built using block encoding have been excluded for the same reason (for example, the discrete adiabatic method [costa2022optimal] and the Zeno eigenstate filtering method [Lin2020optimalpolynomial]). Furthermore, algorithms such as AQC(exp) and CKS approaches that we consider in our study also scale near-optimally in $\kappa$ and $s$, and thus they may provide a reasonable idea of what to expect in this context from the approaches that we have excluded. Lastly, our analysis includes a ‘dream QLS’ whose scaling is more favourable than any of the quantum linear solvers that we do not consider for our survey.
@@ A.4.2  Scope of our numerical study
  • The classical benchmark algorithm: The runtime complexity for the fictitious CLS that we choose is set to be the same as that of the well-known conjugate gradient method purely due to favourable scaling in $N$, $\kappa$, $s$, and $\epsilon^{-1}$, but the CLS itself should not be thought of as possessing the limitations of the conjugate gradient method such as being restricted only for positive definite matrices. In this context, we acknowledge that other important works exist in literature and by choosing a fictitious CLS as a representative benchmark classical solver to compare QLSs against, we do not do justice to the vast body of literature on classical algorithms for linear systems of equations. We defer a comparison of QLSs against leading classical linear solvers to a future study. Having said this, we pick two realistic specialized classical linear solvers for Laplacians as representative examples to qualitatively comment on our solver as a reasonable candidate. To this end, we pick a graph family from our candidate graph families and carry out a comparison:
@@
      The authors in Ref.[koutis2014approaching] discuss an efficient method for solving systems of linear equations arising from a graph Laplacian ($(N\times N)$, symmetric, and diagonally dominant matrix), whose associated graph has $N$ vertices and $M$ edges, and which scales as $\mathcal{\tilde{O}}(M\mathrm{log}^{2}(N)\mathrm{log}(1/\epsilon))$. Our goal is to compare this solver, which we shall term the KMP solver (based on the authors’ initials), against our CLS, and we do so by analyzing a situation where the former excels, thereby stacking the odds against CLS. We set $\epsilon^{-1}=\mathrm{log}(N)$. Noting that a graph that has no isolated vertices needs to have at least N1N-1 edges (best case scenario; the worst case would be $\sim N^{2}$ edges) and also simplifying $\widetilde{O}(f(n))=O(f(n)\log^{k}n)$ by setting $k=1$ (best case scenario), the expression for complexity is $\mathcal{O}(N\mathrm{log}^{2}(N)\mathrm{log}(1/\epsilon)\ \mathrm{log}(N\mathrm{log}^{2}(N)\mathrm{log}(1/\epsilon)))$. For illustration, we now consider the specific case of a graph whose number of vertices grow as $N=2^{n}$, while $\kappa$ and $s$ grow as log($N$). We see that the runtime complexity ratio, $\tilde{R}(n)=t_{\mathrm{CLS}}/t_{\mathrm{HHL}}$, simplifies to $2^{n}\log(n)/n^{5.5}$, which is an exponential advantage. Instead, if we use the KMP algorithm in place of CLS, the runtime ratio is $(2^{n}\mathrm{log}({n})\log[2^{n}n^{2}\log(n)])/n^{5}$. Fig. A.2(c) presents the curves for $\tilde{R}(n)$ versus $n$ for the cases of CLS and the KMP solver. We see that using either our CLS or the KMP solver still yields an exponential advantage with HHL. We also observe that the crossover point of $\tilde{R}(n)=1$ for the CLS we consider is at $n=24$, that is, $N=16$ million, it occurs at an $n$ value of 14, that is, $N=16000$, for KMP the algorithm. Therefore, one can start seeing an advantage at much smaller system sizes with specialized (and realistic) algorithms such as the KMP approach.
@@
      We consider another example, where the authors propose algorithms for solving systems of linear equations that arise in the specific case of graph Laplacians of planar graphs [koutis2007linear, spielman2010algorithms]. Their algorithms scale as $\mathcal{O}(M\mathrm{log}(1/\epsilon))$. Assuming $M\sim N$, $1/\epsilon\sim\mathrm{log}(N)$ and $N\sim 2^{n}$, we obtain $\tilde{R}$ to be $(2^{n}\mathrm{log}({n}))/n^{7}$, which gives an exponential advantage, just as in the case of CLS being used in place of these approaches. Furthermore, these algorithms too reach the crossover point at smaller system sizes.
    Thus, we see that qualitatively, the CLS we consider is a reasonable representative example at least for the purposes of our preliminary study to all of the classical linear solvers. For completeness, we comment on a special case of LSP applied to effective resistance determination where the conjugate gradient method works. Since the input vector, $\vec{b}$ has in it one $1$, one 1-1, and the rest of its elements as 0, the vector is orthogonal to the all-1 vector that lies in the null space of $L$, and thus conjugate gradient suffices in spite of its limitation in being able to handle only positive definite matrices (for example, see Section 3.2 of Ref. [spielman2010algorithms]). In fact, this holds when the condition $\sum_{i}b_{i}=0$ is satisfied.
@@
Appendix A.5 Details of the graph families surveyed: construction and input parameters
In this section, we list all the classes of graph families which we consider for our numerical investigations. We recall that we consider both random and non-random graphs for our study. A random graph is a graph where the existence of an edge is probabilistic. The probability depends on a seed value. For different seed values, we get different probability functions that result in different graphs belonging to the same class. In a non-random graph, the existence of an edge is deterministic. It is unrealistic to consider all the graphs of the same class of random graphs for a numerical investigation. But fixing the seed value makes a random graph unique, which we consider as a representative of the family of random graphs. In addition, other than the seed values, there can be other parameters that influence the structural properties of the graphs. Therefore, for each of the graphs, we present the seed values and parameters that we set for the numerical calculations in this work. We follow NetworkX (version 3.4.2) for the construction of all random and non-random graphs except the hypercube graphs. In the following two sections, we collect the graphs used for our studies on the systems of linear equations generated by the Laplacian matrices and the incidence matrices.
A.5.1  Graphs whose Laplacian matrices are considered in our study
@@ A.5.1.1 Non-Random graphs
    Hypercube graphs: The hypercube graphs are generalizations of cubes in graph theory. The $n$-dimensional hypercube graph $G_{2}^{n}$ has vertex set $V(G_{2}^{n})=\{0,1\}^{\times n}$ which contains $N=2^{n}$ vertices. The Hamming distance $\mathfrak{H}(x,y)$ between two tuples $x=(x_{1},x_{2},\cdots,x_{n})$ and $y=(y_{1},y_{2},\cdots,y_{n})\in\{0,1\}^{\times n}$ is the number of indices $i$ for which $x_{i}\neq y_{i}$. There is an edge between vertices $x$ and $y$ if the Hamming distance, $\mathfrak{H}(x,y)=1$. Note that $x_{i}$ and $y_{i}$ are either 0 or $1$ for $i=1,2,\cdots,n$. The dimension $n$ is the only parameter for this class of graphs. In this article, we consider all the hypercube graphs with number of vertices $N$ ranging from $4$ to $16384$. We present the illustrations of two hypercube graphs in Figs. 2(a).
@@
    Margulis-Gabber-Galil graphs: In this article, we consider the vertex sets of the Margulis-Gabber-Galil graphs as $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$, where $\mathbb{Z}_{n}=\{\overline{0},\overline{1},\overline{2},\cdots,\overline{(n-1)}\}$, $\overline{k}$ is the set of all integers $k$ modulo $n$, and $n$ is a natural number. Each vertex $(x,y)\in\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ is connected to $(x\oplus 2y,y),(x\oplus 2y\oplus 1,y),(x,y\oplus 2x)$ and $(x,y\oplus 2x\oplus 1)$, where $\oplus$ denotes the addition modulo $n$. This construction would lead to self loops and parallel edges. We modify the graph to exclude self loops and parallel edges from this family. We note that the number of vertices in this family grows as $N=n^{2}$. We consider all such possible modified Margulis-Gabber-Galil graphs with $25\leq N\leq 11881$. We depict modified Margulis-Gabber-Galil graphs for $n=3$ and $n=6$ in Fig. 3(a).
@@
    Sudoku graphs [herzberg2007sudoku]: The Sudoku graphs are graphs that consist of $N=n^{4}$ vertices, which are distributed into the cells of an $n^{2}\times n^{2}$ grid. Every cell contains $n^{2}$ vertices. Two distinct vertices are adjacent if they belong to the same row, column, or cell. Thus, every vertex of Sudoku graph is connected to $3n^{2}-2n-1$ vertices. The quantity $d_{max}$ grows as $n^{2}$, which is also reflected in sparsity of the Laplacian matrix. The In our computations, we consider all the Sudoku graphs whose vertices lie in the range of $16\leq N\leq 104976$. We draw Sudoku graphs in Fig. 3(d) for $n=2$ and $n=3$.
@@
      We fix $r=101$ and vary $c$ from $5$ to $197$ to generate a family of graphs whose vertices lie in the range $602\leq N\leq 20196$. Two grid 2d graphs belonging to this family are depicted in Fig. A.9(a).
@@
      For the other family of graphs we consider $r+1=c+1=2^{n}$. With this constraint on the number of vertices, we generate graphs with $16\leq N\leq 65536$. Grid 2d graphs belonging to this family for $n=2$ and $n=3$ are depicted in Fig. A.9(bi).
@@
    Hexagonal lattice graphs: The graph is constructed on a two dimensional plane having $r$ rows and $c$ columns of hexagons. The vertices of the hexagons constitute the vertices of the graph and the edges of the graph are the sides of the hexagons (see Fig. A.9(d)). For our analysis, we set $c=101$ and generate the family of graphs with $r=1$ to $r=108$. Thus, the number of vertices lie in the range $406\leq N\leq 22234$.
@@
    Triangular lattice graphs: The construction involves arranging triangles in $r$ rows and $c$ columns on a two-dimensional plane. The corners of the triangles are the vertices of the graph. Also, the sides of the triangles are the edges of the graph. The number of rows $r$ is varied from $1$ to $100$ while keeping the number of columns $c$ fixed at $101$. The graphs are generated for $N$ in the range $103\leq N\leq 5202$. Two instances of triangular lattice graphs are presented in Fig. A.9(a).
@@
    Complete graphs: A complete graph is a graph, such that an edge between any two distinct vertices exists. The degree of all the vertices is N1N-1 in a complete graph with $N$ vertices. We consider all the complete graphs with number of vertices $2\leq N\leq 5004$. Two instances of this graph can be found in Fig. A.9(j).
@@
    Turan graphs: In general, Turan graphs are constructed from $p$ partitions or disjoint sets of different sizes. For our calculations, we set $p=2$, making it a bipartite graph. The number of vertices in those two partitions are $\big\lfloor\frac{N}{2}\big\rfloor$ and $N-\big\lfloor\frac{N}{2}\big\rfloor$. We analyze graphs whose vertices are in the range $5\leq N\leq 5003$. Two illustrations of the graph family can be found in Fig. A.9(m).
@@
    $H_{k,n}$ Harary graphs [harary1962maximum]: For this family of graphs, we set the vertex connectivity, which refers to the minimum number of edges to be removed from the graph to make it a disconnected graph, to $3$. The number of vertices $N=n$ is varied from $5$ to $5009$. The construction procedure tries to minimize the number of edges in the graph. Two instances of the graph family are presented in Fig. A.9(ak).
@@
    $H_{m,n}$ Harary graphs [harary1962maximum] : Given $m$ edges and $N=n$ vertices, this construction yields a graph that ensures maximum vertex connectivity. The number of edges in this family of graphs is set to be one greater than the number of vertices that the graph contains. Here, we generate graphs for $5\leq N\leq 5009$. We present two graphs from this family in Fig. A.9(an).
@@
    Ladder graphs: A path graph $P=(V,E)$ of order $n$ is a graph with vertex set $V=\{v_{1},v_{2},\cdots,v_{n}\}$ and edge set $E=\{(v_{i},v_{i+1}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,(n-1)\}$. A ladder graph with $N=2n$ vertices is constructed with all the vertices and edges of two path graphs $P_{1}$ and $P_{2}$ each of order $n$ as well as the edges $\{(u_{i},v_{i}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,n\penalty 10000\ \text{and}\penalty 10000\ u_{i}\in V(P_{1}),v_{i}\in V(P_{2})\}$. We consider all the ladder graphs with $10\leq N\leq 5000$. Two graphs of this family are presented in Fig. A.9(at).
@@
    Circular ladder graphs: A cycle graph $C=(V,E)$ of order $n$ is a graph with vertex set $V=\{v_{1},v_{2},\cdots,v_{n}\}$ and edge set $E=\{(v_{i},v_{i+1}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,(n-1)\}\cup\{(v_{n},v_{1})\}$. A circular ladder graph with $N=2n$ vertices is constructed with all the vertices and edges of two cycle graphs $C_{1}$ and $C_{2}$ each of order $n$ as well as the edges $\{(u_{i},v_{i}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,n\penalty 10000\ \text{and}\penalty 10000\ u_{i}\in V(P_{1}),v_{i}\in V(P_{2})\}$. In this case also, we consider all the circular ladder graphs with $10\leq N\leq 5000$. Two graphs of this family are given in Fig. A.9(aq).
@@
    Ring of cliques: A clique is a complete subgraph of a graph. A ring of clique is constructed by connecting each clique by a single edge. In this work, we set the number of vertices in each clique to be $3$, and thus all cliques are of equal size. Hence, the number of vertices in the graph grows as $N=3n$, for $n$ cliques. We consider all possible graphs with vertices in the range $15\leq N\leq 5001$. Fig. A.9(aw).
@@
      Balanced binary tree: In a balanced binary tree the degree of the root vertex $r=2$. If the height of the tree is $n$, then the number of vertices, $N=2^{n+1}-1$. For our calculations, we consider $n$ from $2$ to $14$. Two graphs from the graph family are presented in Fig. A.9(az).
@@
      Balanced ternary tree: Similarly, in a balanced ternary tree the degree of root vertex $r=3$. If the height of the tree is $n$, then the number of vertices, $N=(3^{n+1}-1)/2$. For our calculations, we consider $n$ from $2$ to $9$. Two instances are presented in Fig. A.9(bc).
@@
    Binomial tree: A binomial tree of order $n$ is created by linking the root vertices of two identical copies of binomial trees of order n1n-1. A binomial tree of order 0 has a single vertex and this vertex acts as a root vertex in successive graphs. Suppose that the root vertex of binomial tree of order n1n-1 is $v_{1}$. We create its identical copy, whose root vertex is now marked as $v_{c}$, and join $v_{1}$ and $v_{c}$ to construct a binomial tree of order $n$. The root vertex $v_{c}$ can be viewed as the leftmost child of $v_{1}$. The number of vertices, $N$, in the tree grows as $N=2^{n}$. For our calculations, we consider $n$ from $2$ to $14$. Two instances of the graph family are given in Fig. A.9(bf).
A.5.1.2 Random graphs
@@
    Random regular expander graphs: Our procedure for constructing the random regular expander graph is as follows. We generate a random regular graph with regularity $k$ and $N$ vertices using the construction prescribed in Ref. [friedman2003proof]. Next, we check if the second largest eigenvalue of the adjacency matrix $\lambda_{2}\leq 2\sqrt{k-1}$, where $k$ is an even number. In our construction, we set $k=6$. Recall that all the Ramanujan graphs [murty2020ramanujan], which is a prominent family of expander graphs, must fulfill this property. We repeat the procedure at most $200$ times for getting a Ramanujan graph. We experience that this process could not produce a Ramanujan graph in $3$ percent of the considered values of $N$, which is a negligible percentage. Varying the value of $N$ between $9$ and $5009$ we construct a family of Ramanujan graphs, used in our investigation. We present two graphs of this family in Fig. A.8(g).

### v4

@@ Acknowledgements.
  • Choice of other graph input parameters: We saw earlier the effect of edge weight choice on the prospect of realizing an advantage. Other graph input parameters such as choice of seed value may also influence our conclusions. For our analyses presented in the main text, we fix the seed value for reproducibility of our results. To understand the influence of seed values, we pick three representative candidates: Barabási-Albert graph family from Laplacian matrix based system of linear equations, Gaussian random partition graph (no source and sink) family, and planted partition graph (no source and sink) family from incidence matrix based system of linear equations. The results of our analyses are presented in Fig. S.7. For each graph family, we pick three different seed values and analyze the advantage offered by the respective graph families. We find that for the candidates that we consider, there is no change in their category upon changing the seed values. For example, the Barabási-Albert graph, which is a good graph, remains a good graph. Although the categories do not change for the graph families for the seed values that we considered for them, our observations cannot be generalized, and this aspect requires further analysis in a future study. The effect of tuning other parameters on advantage is also beyond the scope of our current work, and we defer it for a future study.
S.4.2  Scope of our numerical study
  • Number of quantum linear solvers considered: Although we consider several efficient QLSs for our survey, we do exclude some, such as the quantum singular value transformation (QSVT)-based QLS. While QSVT is important due to its potential to offer near-optimal scaling in $\kappa$ and $\epsilon$, we do not consider it due to the non-trivial nature of $\mathcal{N}$ dependence entering via block encoding costs in its complexity. Other near-optimally scaling approaches built using block encoding have been excluded for the same reason (for example, the discrete adiabatic method [17] and the Zeno eigenstate filtering method [40]). Furthermore, algorithms such as AQC(exp) and CKS approaches that we consider in our study also scale near-optimally in $\kappa$ and $s$, and thus they may provide a reasonable idea of what to expect in this context from the approaches that we have excluded. Lastly, our analysis includes a ‘dream QLS’ whose scaling is more favourable than any of the quantum linear solvers that we do not consider for our survey.
@@
  • The classical benchmark algorithm: The runtime complexity for the fictitious CLS that we choose is set to be the same as that of the well-known conjugate gradient method purely due to favourable scaling in $\mathcal{N}$, $\kappa$, $s$, and $\epsilon^{-1}$, but the CLS itself should not be thought of as possessing the limitations of the conjugate gradient method such as being restricted only for positive definite matrices. In this context, we acknowledge that other important works exist in literature and by choosing a fictitious CLS as a representative benchmark classical solver to compare QLSs against, we do not do justice to the vast body of literature on classical algorithms for linear systems of equations. We defer a comparison of QLSs against leading classical linear solvers to a future study. Having said this, we pick two realistic specialized classical linear solvers for Laplacians as representative examples to qualitatively comment on our solver as a reasonable candidate. To this end, we pick a graph family from our candidate graph families and carry out a comparison:
@@
      The authors in Ref.[35] discuss an efficient method for solving systems of linear equations arising from a graph Laplacian ($(N\times N)$, symmetric, and diagonally dominant matrix), whose associated graph has $N$ vertices and $M$ edges, and which scales as $\mathcal{\tilde{O}}(M\mathrm{log}^{2}(N)\mathrm{log}(1/\epsilon))$. Our goal is to compare this solver, which we shall term the KMP solver (based on the authors’ initials), against our CLS, and we do so by analyzing a situation where the former excels, thereby stacking the odds against CLS. We set $\epsilon^{-1}=\mathrm{log}(N)$. Noting that a graph that has no isolated vertices needs to have at least N1N-1 edges (best case scenario; the worst case would be $\sim N^{2}$ edges) and also simplifying $\widetilde{O}(f(n))=O(f(n)\log^{k}n)$ by setting $k=1$ (best case scenario), the expression for complexity is $\mathcal{O}(N\mathrm{log}^{2}(N)\mathrm{log}(1/\epsilon)\ \mathrm{log}(N\mathrm{log}^{2}(N)\mathrm{log}(1/\epsilon)))$. For illustration, we now consider the specific case of a graph for which $\kappa$ and $s$ grow as log($N$). We see that the runtime complexity ratio, $R(N)=t_{\mathrm{CLS}}/t_{\mathrm{HHL}}$, simplifies to $N\log(\log(N))/\log^{5.5}(N)$, which is an exponential advantage. Instead, if we use the KMP algorithm in place of CLS, the runtime ratio is $\frac{(N\log(\log(N))\log[N\mathrm{log}^{2}(N)\log(\log(N))])}{\log^{5}(N)}$. Fig. S.2(c) presents the curves for R(N){R}(N) versus $N$ for the cases of CLS and the KMP solver. We see that using either our CLS or the KMP solver still yields an exponential advantage with HHL.
@@
      We consider another example, where the authors propose algorithms for solving systems of linear equations that arise in the specific case of graph Laplacians of planar graphs [36, 60]. Their algorithms scale as $\mathcal{O}(M\mathrm{log}(1/\epsilon))$. Assuming $M\sim N$, $1/\epsilon\sim\mathrm{log}(N)$, we obtain $R(N)$ to be $(N\mathrm{log}(\mathrm{log}(N))/\mathrm{log}^{7}(N)$, which gives an exponential advantage, just as in the case of CLS being used in place of these approaches. Furthermore, these algorithms too reach the crossover point at smaller system sizes.
    Thus, we see that qualitatively, the CLS we consider is a reasonable representative example at least for the purposes of our preliminary study to all of the classical linear solvers. For completeness, we comment on a special case of LSP applied to effective resistance determination where the conjugate gradient method works. Since the input vector, $\vec{b}$ has in it one $1$, one 1-1, and the rest of its elements as 0, the vector is orthogonal to the all-1 vector that lies in the null space of $L$, and thus conjugate gradient suffices in spite of its limitation in being able to handle only positive definite matrices (for example, see Section 3.2 of Ref. [60]). In fact, this holds when the condition $\sum_{i}b_{i}=0$ is satisfied.
@@
Appendix S.5 Details of the graph families surveyed: construction and input parameters
In this section, we list all the classes of graph families which we consider for our numerical investigations. We recall that we consider both random and non-random graphs for our study. A random graph is a graph where the existence of an edge is probabilistic. The probability depends on a seed value. For different seed values, we get different probability functions that result in different graphs belonging to the same class. In a non-random graph, the existence of an edge is deterministic. It is unrealistic to consider all the graphs of the same class of random graphs for a numerical investigation. But fixing the seed value makes a random graph unique, which we consider as a representative of the family of random graphs. Other than the seed values, there can be other parameters that influence the structural properties of the graphs. Therefore, for each of the graphs, we present the seed values and parameters that we set for the numerical calculations in this work. We follow NetworkX (version 3.4.2) for the construction of all random and non-random graphs except the hypercube graphs. In the following two sections, we describe the graphs used for our studies on the system of linear equations generated by the Laplacian matrices and the incidence matrices.
S.5.1  Graphs whose Laplacian matrices are considered in our study
@@
    Hypercube graphs: The hypercube graphs are generalizations of cubes in graph theory. The $n$-dimensional hypercube graph $G_{2}^{n}$ has vertex set $V(G_{2}^{n})=\{0,1\}^{\times n}$ which contains $N=2^{n}$ vertices. The Hamming distance $\mathfrak{H}(x,y)$ between two tuples $x=(x_{1},x_{2},\cdots,x_{n})$ and $y=(y_{1},y_{2},\cdots,y_{n})\in\{0,1\}^{\times n}$ is the number of indices $i$ for which $x_{i}\neq y_{i}$. There is an edge between vertices $x$ and $y$ if the Hamming distance, $\mathfrak{H}(x,y)=1$. Note that $x_{i}$ and $y_{i}$ are either 0 or $1$ for $i=1,2,\cdots,n$. The dimension $n$ is the only parameter for this class of graphs. In this article, we consider all the hypercube graphs with number of vertices $N$ ranging from $4$ to $16384$. The illustrations of the hypercube graph can be found in Figure 2(a) of the main manuscript.
@@
    Margulis-Gabber-Galil graphs: In this article, we consider the vertex sets of the Margulis-Gabber-Galil graphs as $\mathbb{Z}_{n}\times\mathbb{Z}_{n}$, where $\mathbb{Z}_{n}=\{\overline{0},\overline{1},\overline{2},\cdots,\overline{(n-1)}\}$, $\overline{k}$ is the set of all integers $k$ modulo $n$, and $n$ is a natural number. Each vertex $(x,y)\in\mathbb{Z}_{n}\times\mathbb{Z}_{n}$ is connected to $(x\oplus 2y,y),(x\oplus 2y\oplus 1,y),(x,y\oplus 2x)$ and $(x,y\oplus 2x\oplus 1)$, where $\oplus$ denotes the addition modulo $n$. This construction would lead to self loops and parallel edges. We modify the graph to exclude self loops and parallel edges from this family. We note that the number of vertices in this family grows as $N=n^{2}$. We consider all such possible modified Margulis-Gabber-Galil graphs with $25\leq N\leq 11881$. We depict modified Margulis-Gabber-Galil graphs for $n=3$ and $n=6$ in Figure 2(c) of the main manuscript.
@@
    Sudoku graphs [32]: The Sudoku graphs are graphs that consist of $N=n^{4}$ vertices, which are distributed into the cells of an $n^{2}\times n^{2}$ grid. Every cell contains $n^{2}$ vertices. Two distinct vertices are adjacent if they belong to the same row, column, or cell. Thus, every vertex of Sudoku graph is connected to $3n^{2}-2n-1$ vertices. The quantity $d_{max}$ grows as $n^{2}$, which is also reflected in sparsity of the Laplacian matrix. In our computations, we consider all the Sudoku graphs whose vertices lie in the range of $16\leq N\leq 104976$. We draw Sudoku graphs in Fig. S.8(a) for $n=2$ and $n=3$.
@@
      We fix $r=101$ and vary $c$ from $5$ to $197$ to generate a family of graphs whose vertices lie in the range $602\leq N\leq 20196$. Two grid 2d graphs belonging to this family are depicted in Figure 3(a) of the main manuscript.
@@
      For the other family of graphs we consider $r+1=c+1=2^{n}$. With this constraint on the number of vertices, we generate graphs with $16\leq N\leq 65536$. Grid 2d graphs belonging to this family for $n=2$ and $n=3$ are depicted in Fig. S.9(ak).
@@
    Hexagonal lattice graphs: The graph is constructed on a two dimensional plane having $r$ rows and $c$ columns of hexagons. The vertices of the hexagons constitute the vertices of the graph and the edges of the graph are the sides of the hexagons (see Figure 3(c) of the main manuscript). For our analysis, we set $c=101$ and generate the family of graphs with $r=1$ to $r=108$. Thus, the number of vertices lie in the range $406\leq N\leq 22234$.
@@
    Triangular lattice graphs: The construction involves arranging triangles in $r$ rows and $c$ columns on a two-dimensional plane. The corners of the triangles are the vertices of the graph. Also, the sides of the triangles are the edges of the graph. The number of rows $r$ is varied from $1$ to $100$ while keeping the number of columns $c$ fixed at $101$. The graphs are generated for $N$ in the range $103\leq N\leq 5202$. Two instances of triangular lattice graphs are presented in Fig. S.9(a).
@@
    Complete graphs: A complete graph is a graph, such that an edge between any two distinct vertices exists. The degree of all the vertices is N1N-1 in a complete graph with $N$ vertices. We consider all the complete graphs with number of vertices $2\leq N\leq 5004$. Two instances of this graph can be found in Fig. S.9(c).
@@
    Turan graphs: In general, Turan graphs are constructed from $p$ partitions or disjoint sets of different sizes. For our calculations, we set $p=2$, making it a bipartite graph. The number of vertices in those two partitions are $\big\lfloor\frac{N}{2}\big\rfloor$ and $N-\big\lfloor\frac{N}{2}\big\rfloor$. We analyze graphs whose vertices are in the range $5\leq N\leq 5003$. Two illustrations of the graph family can be found in Fig. S.9(e).
@@
    $H_{k,n}$ Harary graphs [30]: For this family of graphs, we set the vertex connectivity, which refers to the minimum number of edges to be removed from the graph to make it a disconnected graph, to $3$. The number of vertices $N=n$ is varied from $5$ to $5009$. The construction procedure tries to minimize the number of edges in the graph. Two instances of the graph family are presented in Fig. S.9(u).
@@
    $H_{m,n}$ Harary graphs [30] : Given $m$ edges and $N=n$ vertices, this construction yields a graph that ensures maximum vertex connectivity. The number of edges in this family of graphs is set to be one greater than the number of vertices that the graph contains. Here, we generate graphs for $5\leq N\leq 5009$. We present two graphs from this family in Fig. S.9(w).
@@
    Ladder graphs: A path graph $P=(V,E)$ of order $n$ is a graph with vertex set $V=\{v_{1},v_{2},\cdots,v_{n}\}$ and edge set $E=\{(v_{i},v_{i+1}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,(n-1)\}$. A ladder graph with $N=2n$ vertices is constructed with all the vertices and edges of two path graphs $P_{1}$ and $P_{2}$ each of order $n$ as well as the edges $\{(u_{i},v_{i}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,n\penalty 10000\ \text{and}\penalty 10000\ u_{i}\in V(P_{1}),v_{i}\in V(P_{2})\}$. We consider all the ladder graphs with $10\leq N\leq 5000$. Two graphs of this family are presented in Fig. S.9(aa).
@@
    Circular ladder graphs: A cycle graph $C=(V,E)$ of order $n$ is a graph with vertex set $V=\{v_{1},v_{2},\cdots,v_{n}\}$ and edge set $E=\{(v_{i},v_{i+1}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,(n-1)\}\cup\{(v_{n},v_{1})\}$. A circular ladder graph with $N=2n$ vertices is constructed with all the vertices and edges of two cycle graphs $C_{1}$ and $C_{2}$ each of order $n$ as well as the edges $\{(u_{i},v_{i}):\penalty 10000\ \text{for}\penalty 10000\ i=1,2,\cdots,n\penalty 10000\ \text{and}\penalty 10000\ u_{i}\in V(P_{1}),v_{i}\in V(P_{2})\}$. In this case also, we consider all the circular ladder graphs with $10\leq N\leq 5000$. Two graphs of this family are given in Fig. S.9(y).
@@
    Ring of cliques: A clique is a complete subgraph of a graph. A ring of clique is constructed by connecting each clique by a single edge. In this work, we set the number of vertices in each clique to be $3$, and thus all cliques are of equal size. Hence, the number of vertices in the graph grows as $N=3n$, for $n$ cliques. We consider all possible graphs with vertices in the range $15\leq N\leq 5001$. Fig. S.9(ac).
@@
      Balanced binary tree: In a balanced binary tree the degree of the root vertex $r=2$. If the height of the tree is $n$, then the number of vertices, $N=2^{n+1}-1$. For our calculations, we consider $n$ from $2$ to $14$. Two graphs from the graph family are presented in Fig. S.9(ae).
@@
      Balanced ternary tree: Similarly, in a balanced ternary tree the degree of root vertex $r=3$. If the height of the tree is $n$, then the number of vertices, $N=(3^{n+1}-1)/2$. For our calculations, we consider $n$ from $2$ to $9$. Two instances are presented in Fig. S.9(ag).
@@
    Binomial tree: A binomial tree of order $n$ is created by linking the root vertices of two identical copies of binomial trees of order n1n-1. A binomial tree of order 0 has a single vertex and this vertex acts as a root vertex in successive graphs. Suppose that the root vertex of binomial tree of order n1n-1 is $v_{1}$. We create its identical copy, whose root vertex is now marked as $v_{c}$, and join $v_{1}$ and $v_{c}$ to construct a binomial tree of order $n$. The root vertex $v_{c}$ can be viewed as the leftmost child of $v_{1}$. The number of vertices, $N$, in the tree grows as $N=2^{n}$. For our calculations, we consider $n$ from $2$ to $14$. Two instances of the graph family are given in Fig. S.9(ai).
S.5.1.2 Random graphs
@@
    Random regular expander graphs: Our procedure for constructing the random regular expander graph is as follows. We generate a random regular graph with regularity $k$ and $N$ vertices using the construction prescribed in Ref. [23]. Next, we check if the second largest eigenvalue of the adjacency matrix $\lambda_{2}\leq 2\sqrt{k-1}$, where $k$ is an even number. In our construction, we set $k=6$. Recall that all the Ramanujan graphs [48], which is a prominent family of expander graphs, must fulfill this property. We repeat the procedure at most $200$ times for getting a Ramanujan graph. We experience that this process could not produce a Ramanujan graph in $3$ percent of the considered values of $N$, which is a negligible percentage. Varying the value of $N$ between $9$ and $5009$ we construct a family of Ramanujan graphs, used in our investigation. We present two graphs of this family in Fig. S.8(g).